OpenStudy (kamille):

Hello. I need some help There is a code, which is made by two letters from 23 letters and 3 numbers from numbers 2,3,4. This code doesnt exist, when all three numbers in a code are the same. How many ways a code can be made. How I think: The letters can be: 23*23=529 Numbers: 3*3*2=18 But the answer is 12144. Can anyone help me?

OpenStudy (kamille):

Omg! I have forgotten to say that there is made of 5signs

OpenStudy (anonymous):

i thought we have "26" alphabets

OpenStudy (nathan917):

Yeah

OpenStudy (kamille):

Well, it is how the problem sounds. We can take any number, I need a way how to solve this problem

OpenStudy (kamille):

and, actually, I am NOT from USA/UK and our language doesnt have Q,W,X so it is why 23:))

OpenStudy (kamille):

so, any tips?

OpenStudy (anonymous):

two letters from 23 = 23*22 3 letters = 3*2*1 = 6

OpenStudy (anonymous):

how many signs you need in the code?

OpenStudy (kamille):

5

OpenStudy (anonymous):

"5" available and you use how many of them?

OpenStudy (kamille):

and why it is 3*2*1 if it can be two same numbers?

OpenStudy (kamille):

all of them

OpenStudy (anonymous):

yes. the order.. you can arrange them in 6 ways

OpenStudy (anonymous):

impossible.. all "5" signs? your answer uses one of "4" signs

OpenStudy (anonymous):

if "5" is correct, then the answer will not end with a "4" at its units place.

OpenStudy (kamille):

why it is not 3*3*2? :) You can use TWO same numbers, but you cant use three same numbers. For example: You CAN use a code like this AB 233, but you CANT use a code AB 333

OpenStudy (anonymous):

could you post the exact wordings of the question.

OpenStudy (anonymous):

so, it is then alphabets = 23*22 numbers = 3^3-3 = 24 total combinations = 23*22*24

OpenStudy (kamille):

Can you explain me?

OpenStudy (anonymous):

alphabets = no repetition = 23*22 kapeesh?

OpenStudy (anonymous):

number of ways of arranging all "3" numbers = 3^3 but no repetition -> so -> 3^3-3

OpenStudy (kamille):

well, the problem doesnt say letters cannot repeat:) But the answer is correct, thank you.

OpenStudy (anonymous):

I know. and the signs are unacconted for too. but this is the only way I can explain that number you gave as the answer!

OpenStudy (kamille):

oh,okay:) Thank you.

OpenStudy (anonymous):

yer welcome :D